### v1

@@ 1 Introduction
Merger evaluation remains one of the most challenging problems in antitrust policy because of the difficulties separating efficiency gains from increased market power. This literature has developed two distinct methodologies: reduced form retrospective analyses that examine actual post-merger outcomes by applying event study methods like difference-in-differences (DiD) to pre- and post-merger data (e.g., Kim and Singal, 1993; Focarelli and Panetta, 2003; Hastings, 2004; Ashenfelter and Hosken, 2010; Ashenfelter et al., 2015; Carlton et al., 2019), and structural prospective analyses that predict merger effects using only pre-merger data (e.g., Nevo, 2000; Nevo and Whinston, 2010). There are a few important exceptions to this classification (Baker, 1999; Pesendorfer, 2003; Davis, 2005; Miller and Weinberg, 2017; Igami and Sugaya, 2022; Bruegge et al., 2024), which underpin our work here. For a more comprehensive overview of mergers see Whinston (2006) Chapter 3, Kwoka, Jr. (2014) and Asker and Nocke (2021).
@@
Our paper follows a systematic analysis that sequentially integrates reduced-form and structural approaches. We begin with a reduced-form retrospective analysis of pre- and post-merger data (Section 3). We then evaluate the robustness of this approach concerning the time trends to capture possibly declining demand. Next, we develop a structural interpretation of these reduced-form estimates, explicitly delineating the economic assumptions underlying the DiD strategy (Section 4), and find that the structural estimates suggest mergers generate cost efficiency and lead to a decline in competitiveness. Then, we develop a reduced-form approach that relies on the exogenous variation in market structure in only the pre-merger data to forecast merger effects (Section 5). We conclude by implementing a structural merger simulation that employs only the pre-merger data and uses estimated effects of mergers on prices from our Sections 3 and 5 to determine the post-merger marginal costs and use it to evaluate the reduced form estimates under different post-merger strategic conduct based on the findings in Section 4 (Section 6).
@@
Our structural model reveals three critical insights about the model implicitly assumed in a typical retrospective analysis. First, consistent DiD estimation requires that mergers not alter firm conduct and that time-varying unobservable “demand shocks” evolve similarly across treated and control markets. Second, the error term in the DiD specification contains unobservable demand and supply shocks. While the cost shocks may reasonably be uncorrelated with merger decisions, demand shocks could systematically differ between treated and control markets—for instance, if firms are more likely to merge in markets with declining demand—and biasing the DiD estimates. Third, our estimates suggest that mergers affect firm conduct differently across market structures, indicating both unilateral and coordinated effects that the retrospective analysis might miss.
@@
Next, we consider the scenario typically faced by the antitrust authorities and the courts: evaluating mergers using only pre-merger data. In Section 5, we develop a reduced-form approach that leverages exogenous variations in market structure pre-merger to predict post-merger effects using only pre-merger data. Building on methodology pioneered by Baker (1999), this approach analyzes how prices vary with different market structures in the pre-merger period to forecast merger-induced changes. Using the first differences method, we find that transitions from duopoly to monopoly are associated with modest but consistent fare increases (ranging from 2.9% to 6.2%) and substantial capacity reductions (15.7% to 27.6%). Conversely, markets transitioning from monopoly to duopoly show symmetric effects: fare decreases of 3.0% to 7.7%, and large capacity increases of 31.5% to 38.5%.
@@
Finally, in Section 6, we implement a structural merger simulation (Nevo, 2000) that considers different scenarios for post-merger cost efficiencies and conduct. In particular, we consider a random utility discrete choice model of demand with differentiated products and Bertrand competition where firms choose prices. Following Verboven (1996), we employ a nested logit demand model to distinguish between nonstop and connecting services, allowing for rich substitution patterns across products and carriers. Our merger simulation analysis reveals that predicted price effects depend critically on assumptions about post-merger cost efficiencies and firms’ conduct. In particular, we explore how post-merger prices depend on efficiency gains and changes in competitive conduct. Under standard competition assumptions and maximum efficiency gains (where merged entities adopt the lower costs of merging firms), the model predicts price decreases ranging from 2.9% to 5.1%. However, this finding presents a puzzle when compared to our earlier findings of stable post-merger prices: either efficiency gains were limited or significant efficiencies were offset by an increase in coordinated effects.
@@
Thus, our bridging framework reveals a crucial tension in evaluating airline mergers that traditional approaches miss. As post-merger prices remain stable, there are two possible interpretations: either the merger generated substantial efficiency gains but also increased market power significantly, or both efficiency gains and competitive effects were modest. This tradeoff creates a strategic dilemma for merging airlines - they can either claim large efficiency gains but must also accept the evidence of increased market power or argue for modest competitive effects. However, then it weakens the case for the merger based on efficiency. By iterating between reduced-form and structural approaches, our framework makes this tradeoff explicit, providing authorities with a more complete toolkit for merger evaluation. Moreover, our finding that market structure changes primarily affect capacity rather than prices suggests that antitrust authorities should broaden their focus beyond price effects alone.
@@
Our paper’s contributions to the literature are both methodological and empirical. Building on the extensive merger analysis literature cited throughout this introduction and the rest of the paper, we integrate existing reduced-form retrospective and structural prospective analyses, drawing insights from each while addressing their limitations. Specifically, we demonstrate how retrospective studies that use post-merger data to measure price effects (e.g., Ashenfelter and Hosken, 2010) and prospective approaches that use pre-merger data and focus on structural estimation of efficiency gains (e.g., Farrell and Shapiro, 1990), can be integrated to provide more robust merger analysis, under the assumption of exogenous mergers and market structure. [^2] This assumption is nearly universal in the empirical merger literature, underlying both retrospective analyses and structural merger simulations, with notable exceptions being Dafny (2009); Aguirregabiria and Ho (2012); Ciliberto et al. (2021); Li et al. (2022), and Bontemps et al. (2023) among others.
@@ 4 Structural Interpretation of Retrospective Analysis
Statistical tests strongly reject the null hypothesis of no change in conduct for all three mergers, confirming significant increases in market power. The magnitudes of the conduct parameter changes (ranging from 0.023 to 0.188) are smaller than the 0.37 estimate reported by Miller and Weinberg (2017) for the beer industry after a joint venture. This difference likely reflects the airline industry’s network structure, where the presence of connecting routes and alternative carriers maintains some competitive pressure even after consolidation. However, mergers need not always increase coordinated effects. Igami and Sugaya (2022) demonstrate that mergers can simultaneously reduce competitiveness (unilateral effects) while decreasing collusion (coordinated effects). These findings suggest that coordinated effects depend on specific industry characteristics, firm asymmetries, and the merger itself.
@@ 5 A Reduced-Form Approach to Merger Prediction
Panel A of Table 6 presents the estimated coefficients ($\Psi$) for the effect of $\Delta HHI$ on prices, passenger volume, and offered seats. The results show a consistent relationship between increased market concentration and higher prices across all three mergers. A one-unit increase in HHI (measured on a 0-1 scale) is associated with price increases ranging from 2.1% (UA-CO) to 10.2% (DL-NW). The DL-NW and AA-US mergers also significantly negatively affect capacity (seats), suggesting that concentrated markets experience higher prices and reduced service. These findings validate our approach of using natural transitions in market structure to forecast merger effects.
@@
Our reduced-form analysis of natural market structure transitions reveals consistent patterns across all three airline mergers. Consolidation leads to moderate price increases but substantial capacity reductions, while market entry produces symmetric effects in the opposite direction. In the next section, we consider a structural approach to prospective merger analysis that incorporates the lessons from these natural experiments and the efficiency considerations from our retrospective evaluation. Our findings differ from Bruegge et al. (2024) in their investigation of the Jet Blue and Spirit Air merger, where they observed that ”entries with larger capacity tend to have greater fare effects, and that the difference in fare effects between carriers varies with the capacity level.” In contrast, our analysis shows significant capacity changes but relatively modest price effects across the mergers we studied.
@@ 7 Conclusion
Looking ahead, promising directions for future research include extending our framework to incorporate endogenous market structure (e.g., Ciliberto et al., 2021), product characteristics and allowing these to change with mergers (e.g., Sweeting, 2010; Fan, 2013; Wollmann, 2018; Li et al., 2022; Bontemps et al., 2023). The challenge will be maintaining our current approach’s transparency while adding these additional features. Our methodology of connecting reduced-form and structural approaches provides a valuable template for such extensions.
@@ B.6 Synthetic DiD
We apply this approach using the implementation package developed by Clarke et al. (2023) to reassess the findings in Table 3. The estimates are presented in Table B.5. These effects suggest three key results. First, we find limited evidence of merger effects on prices and passenger volumes. For instance, the UA-CO merger shows a minimal fare effect of -0.4% in levels and -0.8% in first differences (the latter being marginally significant). The passenger volume effects are similarly modest, approximately 1% in both specifications, and statistically indistinguishable from zero. These estimates suggest economically negligible price impacts from the merger.
@@ C Demand Estimation
Finally, IV Sets 6 and 7 adopt the traditional approach to instruments in the spirit of Bresnahan (1989) and Berry et al. (1995), using the sum of competitors’ origin networks as instruments. Set 6 treats nonstop service as exogenous, while Set 7 treats it as endogenous. Remarkably, both specifications produce results that are broadly consistent with our main findings. Set 6 yields a price coefficient of -1.189 and an elasticity of -3.069, while Set 7 generates similar estimates with a price coefficient of -1.176 and an elasticity of -3.123. The nesting parameters in both specifications satisfy theoretical restrictions (Set 6: $\sigma_{1}$ = 0.153, $\sigma_{2}$ = 0.067; Set 7: $\sigma_{1}$ = 0.191, $\sigma_{2}$ = 0.034) and remain in plausible ranges.
@@
Note: The left panel shows price coefficients and the right panel displays the kernel density estimates of implied unobserved product characteristics $\xi_{j}$ from demand estimation for all three mergers (DL-NW, UA-CO, and AA-US). For each merger, we estimate demand using seven specifications: (1-2) markets connecting top 50 MSAs three years pre-merger; (3) markets connecting top 50 MSAs three years post-merger; (4) markets connecting top 50 MSAs one-year post-merger; (5) three-year post-merger sample with exogenous variables from three years pre-merger; (6) one-year post-merger sample with exogenous variables from one-year pre-merger; and (7) three-year post-merger sample with exogenous variables from one-year pre-merger. These specifications allow us to examine the stability of demand parameters across different time horizons and test the exogeneity of our instruments.

### v2

@@ 1 Introduction
Merger evaluation remains one of the most challenging problems in antitrust policy because of the difficulties separating efficiency gains from increased market power. This literature has developed two distinct methodologies: retrospective analyses that examine post-merger outcomes by applying event study methods like difference-in-differences (DiD) to pre- and post-merger data (e.g., Kim and Singal, 1993; Focarelli and Panetta, 2003; Hastings, 2004; Ashenfelter and Hosken, 2010; Ashenfelter et al., 2015; Carlton et al., 2019), and prospective analyses that predict merger effects using only pre-merger data (e.g., Nevo, 2000; Nevo and Whinston, 2010). There are a few important exceptions to this classification (Baker, 1999; Pesendorfer, 2003; Davis, 2005; Miller and Weinberg, 2017; Igami and Sugaya, 2022; Bruegge et al., 2024), which underpin our work. For a more comprehensive overview of mergers see Whinston (2006); Kwoka, Jr. (2014); Asker and Nocke (2021) and Kaplow (2024).
@@
Our paper follows a systematic analysis that sequentially integrates reduced-form and structural approaches. We begin with a reduced-form retrospective analysis of pre- and post-merger data (Section 3), then evaluate the robustness of this approach concerning the time trends to capture possibly declining demand. Next, we develop an equilibrium model with linear market-level demand that links pricing behavior to mergers (Section 4). This model allows us to interpret the DiD estimation equation from Section 3 as a reduced-form expression of our structural pricing equation, explicitly delineating the economic assumptions that underpin the DiD strategy. Then, we develop a regression-based approach that relies on the exogenous variation in market structure in only the pre-merger data to forecast merger effects (Section 5). We conclude by implementing a structural merger simulation that employs only the pre-merger data and uses estimated effects of mergers on prices from our Sections 3 and 5 to determine the post-merger marginal costs and use it to evaluate the reduced form estimates under different post-merger strategic conduct based on the findings in Section 4 (Section 6).
@@
Our equilibrium model uncovers three critical insights in the implicit assumptions of standard retrospective analyses. First, consistent DiD estimation requires that mergers not alter firm conduct and that time-varying unobservable “demand shocks” evolve similarly across treated and control markets. Second, the error term in the DiD specification contains unobservable demand and supply shocks. While the cost shocks may reasonably be uncorrelated with merger decisions, demand shocks could systematically differ between treated and control markets—for instance, if firms are more likely to merge in markets with declining demand—and biasing the DiD estimates. Third, our estimates suggest that mergers affect firm conduct differently across market structures, indicating both unilateral and coordinated effects that the retrospective analysis might miss.
@@
Next, we consider the scenario typically faced by the antitrust authorities and the courts: evaluating mergers using only pre-merger data. In Section 5, we develop a regression-based approach that leverages exogenous variations in market structure pre-merger to predict post-merger effects using only pre-merger data. Building on methodology pioneered by Baker (1999), this approach analyzes how prices vary with different market structures in the pre-merger period to forecast merger-induced changes. The variation in market structure offers insights into how similar structural changes might affect post-merger outcomes. We find that transitions from duopoly to monopoly are associated with modest but consistent fare increases (ranging from 2.9% to 6.2%) and substantial capacity reductions (15.7% to 27.6%). Conversely, markets transitioning from monopoly to duopoly show symmetric effects: fare decreases of 3.0% to 7.7%, and large capacity increases of 31.5% to 38.5%. While intuitive and simple, this model also captures efficiency induced by past changes in market structure. When a market evolved from a triopoly to a duopoly, the remaining airlines likely optimized their operations in response—and those efficiencies are present in the observed price changes.
@@
Finally, in Section 6, we implement an equilibrium-based merger simulation (Nevo, 2000) that considers different scenarios for post-merger cost efficiencies and conduct. In particular, we consider a random utility discrete choice model of demand with differentiated products and Bertrand competition where firms choose prices. Following Verboven (1996), we employ a nested logit demand model to distinguish between nonstop and connecting services, allowing for rich substitution patterns across products and carriers. Our merger simulation analysis reveals that predicted price effects depend critically on assumptions about post-merger cost efficiencies and firms’ conduct. In particular, we explore how post-merger prices depend on efficiency gains and changes in competitive conduct. Under standard competition assumptions and maximum efficiency gains (where merged entities adopt the lower costs of merging firms), the model predicts price decreases ranging from 2.9% to 5.1%. However, this finding presents a puzzle when compared to our earlier findings of stable post-merger prices: either efficiency gains were limited or significant efficiencies were offset by an increase in coordinated effects.
@@
Thus, our bridging framework reveals a crucial tension in evaluating airline mergers that traditional approaches miss. As post-merger prices remain stable, there are two possible interpretations: either the merger generated substantial efficiency gains but also increased market power significantly, or both efficiency gains and competitive effects were modest. This tradeoff creates a strategic dilemma for merging airlines - they can either claim large efficiency gains but must also accept the evidence of increased market power or argue for modest competitive effects. However, then it weakens the case for the merger based on efficiency. By iterating between reduced-form and structural approaches, our framework makes this tradeoff explicit, providing authorities with a more complete toolkit for merger evaluation. Moreover, our finding that mergers primarily affect capacity rather than prices suggests that antitrust authorities should broaden their focus beyond price effects alone.
@@
Our paper’s contributions to the literature are both methodological and empirical. Building on the extensive merger analysis literature cited throughout this introduction and the rest of the paper, we integrate existing reduced-form retrospective and structural prospective analyses, drawing insights from each while addressing their limitations. Specifically, we demonstrate how retrospective studies that use post-merger data to measure price effects (Ashenfelter and Hosken, 2010) and prospective approaches that use pre-merger data and focus on structural estimation of efficiency gains (Farrell and Shapiro, 1990) can be integrated to provide more robust merger analysis, under the assumption of exogenous mergers. [^2] This assumption is nearly universal in the empirical merger literature, underlying both retrospective analyses and structural merger simulations, with notable exceptions being Dafny (2009); Aguirregabiria and Ho (2012); Ciliberto et al. (2021); Li et al. (2022), and Bontemps et al. (2023) among others.
@@ 4 Structural Interpretation of Retrospective Analysis
Statistical tests strongly reject the null hypothesis of no change in conduct for all three mergers, confirming significant increases in market power. The magnitudes of the conduct parameter changes (ranging from 0.023 to 0.188) are smaller than the 0.37 estimate reported by Miller and Weinberg (2017) for the beer industry after a joint venture. This difference likely reflects the airline industry’s network structure, where the presence of connecting routes and alternative carriers maintains some competitive pressure even after consolidation. However, mergers need not always increase coordinated effects. Igami and Sugaya (2022) demonstrate that mergers can simultaneously reduce competitiveness (unilateral effects) while decreasing collusion (coordinated effects). These findings suggest that coordinated effects depend on specific industry characteristics, firm asymmetries, and mergers.
@@ 5 A Reduced-Form Approach to Merger Prediction
Panel A of Table 6 presents the estimated coefficients ($\Psi$) for the effect of $\Delta HHI$ on prices, passenger volume, and offered seats for the Top 50 markets (by population). The results show a consistent relationship between increased market concentration and higher prices across all three mergers. A one-unit increase in HHI (measured on a 0-1 scale) is associated with price increases ranging from 2.1% (UA-CO) to 10.2% (DL-NW). The DL-NW and AA-US mergers also significantly negatively affect capacity (seats), suggesting that concentrated markets experience higher prices and reduced service. These findings validate our approach of using natural transitions in market structure to forecast merger effects.
@@
Our analysis of exogenous market structure transitions reveals consistent patterns across all three airline mergers. Consolidation leads to moderate price increases but substantial capacity reductions, while market entry produces symmetric effects in the opposite direction. Bruegge et al. (2024), in their investigation of the Jet Blue and Spirit Air merger, where they observed that “entries with larger capacity tend to have greater fare effects, and that the difference in fare effects between carriers varies with the capacity level.” Our analysis shows significant capacity changes but relatively modest price effects across the mergers we studied.
Even though we only use pre-merger data, our methodology captures one of the two components of merger-induced efficiencies: those arising from market structure changes but not those unique to specific mergers. When markets transition from duopoly to monopoly, the efficiency gains achieved by remaining firms are reflected in observed price changes. For instance, when US Airways exits a route previously served alongside American Airlines, the resulting monopoly likely drives American to adjust its operational approach—effects our analysis captures and is informative about outcomes post AA-US merger. Our approach may not fully account for merger-specific efficiencies, such as proprietary fleet optimization between two specific carriers. Nevertheless, this exercise provides valuable real-world evidence of competitive dynamics. In the next section, we consider a structural approach to prospective merger analysis that incorporates the lessons from these natural experiments and the efficiency considerations from our retrospective evaluation.
@@ 7 Conclusion
Looking ahead, promising directions for future research include extending our framework to incorporate endogenous market structure (e.g., Ciliberto et al., 2021; Fan and Yang, Forthcoming), product characteristics and allowing these to change with mergers (e.g., Sweeting, 2010; Fan, 2013; Wollmann, 2018; Li et al., 2022; Bontemps et al., 2023; Atalay et al., 2024). The challenge will be maintaining our current approach’s transparency while adding these additional features. Our methodology of connecting reduced-form and structural approaches provides a valuable template for such extensions.
@@ B.6 Synthetic DiD
We apply this approach using the implementation package developed by Clarke et al. (2023) to reassess the findings in Table 3. The estimates are presented in Table B.5. These effects suggest three key results. First, we find limited evidence of merger effects on prices and passenger volumes. For instance, the UA-CO merger shows a minimal fare effect of -0.4% in levels and -0.8% in first differences (the latter being marginally significant). The passenger volume effects are similarly modest, approximately 1% in both specifications and statistically indistinguishable from zero. These estimates suggest economically negligible price impacts from the merger.
@@ C Demand Estimation
Finally, IV Sets 6 and 7 adopt the traditional approach to instruments in the spirit of Bresnahan (1989) and Berry et al. (1995), using the sum of competitors’ origin networks as instruments. Set 6 treats nonstop service as exogenous, while Set 7 treats it as endogenous. Remarkably, both specifications produce results consistent with our main findings. Set 6 yields a price coefficient of -1.189 and an elasticity of -3.069, while Set 7 generates similar estimates with a price coefficient of -1.176 and an elasticity of -3.123. The nesting parameters in both specifications satisfy theoretical restrictions (Set 6: $\sigma_{1}$ = 0.153, $\sigma_{2}$ = 0.067; Set 7: $\sigma_{1}$ = 0.191, $\sigma_{2}$ = 0.034) and remain in plausible ranges.
@@
Note: The left panel shows price coefficients, and the right panel displays the kernel density estimates of implied unobserved product characteristics $\xi_{j}$ from demand estimation for all three mergers (DL-NW, UA-CO, and AA-US). For each merger, we estimate demand using seven specifications: (1-2) markets connecting top 50 MSAs three years pre-merger; (3) markets connecting top 50 MSAs three years post-merger; (4) markets connecting top 50 MSAs one year post-merger; (5) three-year post-merger sample with exogenous variables from three years pre-merger; (6) one-year post-merger sample with exogenous variables from one-year pre-merger; and (7) three-year post-merger sample with exogenous variables from one-year pre-merger. These specifications allow us to examine the stability of demand parameters across different time horizons and test the exogeneity of our instruments.